OpenStudy (anonymous):

Please Help!!!!!!!!!!!!! Christina keeps her DVD collection in a plastic tub. The table shows the number of each type of DVD in Christina's collection. Christina's DVD Collection tape how many Sci-fi 2 Action 6 Drama 8 Comedy 5 Romance 9 If two DVDs are chosen randomly one at a time, where the first DVD is replaced before the second DVD is chosen, what is the theoretical probability that both DVDs chosen are from the action category? A. 1/25

OpenStudy (anonymous):

B. 1/29 C. 1/30 D. 6/145

OpenStudy (mertsj):

\[\frac{6}{30}\times\frac{6}{30}=\frac{1}{5}\times\frac{1}{5}=\frac{1}{25}\]